OpenStudy (anonymous):

what is the y intercept for 4y+4x=28

OpenStudy (anonymous):

convert to y-intercept form: \(y=mx+b\)\[4y+4x=28\]\[4y=-4x+28\]\[y=-x+7\]the \(b\) in the formula \(y=mx+b\) is the y-intercept...so? :)

OpenStudy (anonymous):

7!? lol

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

4y = 4x - 28 So your y-intercept would be (0,4).

OpenStudy (anonymous):

(0,28) my bad!

OpenStudy (anonymous):

no its not since the entire thing can be divided by 4...it has to be in y-intercept form which means that the y must have a coefficient of 1 and 1 only...its \((0,7)\) buddy :)